0 0
Add a comment Improve this question Transcribed image text
Answer #1
Concepts and reason

The concept used to solve this problem is based on the equilibrium constant of a chemical reaction.

The value of the reaction quotient when a reaction reaches equilibrium is called equilibrium constant. It is the ratio of the concentrations of products and concentrations of reactants.

Fundamentals

For a chemical reaction,

aA+bBcC+dDa{\rm{A}} + b{\rm{B}} \to c{\rm{C}} + d{\rm{D}}

The equilibrium constant is written as follow.

K=[C]c[D]d[A]a[B]bK = \frac{{{{\left[ {\rm{C}} \right]}^c}{{\left[ {\rm{D}} \right]}^d}}}{{{{\left[ {\rm{A}} \right]}^a}{{\left[ {\rm{B}} \right]}^b}}}

Here [C]\left[ {\rm{C}} \right] and [D]\left[ {\rm{D}} \right] are concentrations of products. The concentrations of reactants are [A]\left[ {\rm{A}} \right] and[B]\left[ {\rm{B}} \right].

For the reaction,

The equilibrium constant is written as follow.

K=[B][A]K = \frac{{\left[ {\rm{B}} \right]}}{{\left[ {\rm{A}} \right]}}

The values of K, 0.2 and 9×1099 \times {10^{ - 9}} are less than one. Thus, at these values the reactant A is more than product B at equilibrium.

The values of K, 9000 and 9×1099 \times {10^9} are more than one. Thus, at these values the product B is more than reactant A at equilibrium.

Ans:

The values of K at which B is more than A are 9000 and 9×1099 \times {10^9}.